# Scalable Two-Stage Stochastic Optimal Power Flow via Separable Approximation

### v1

@@ Abstract
Index Terms:
Decision-making under uncertainty, power systems planning, stochastic optimal power flow, value function approximation.
@@ I Introduction
The reliable, secure, and economical operation of power systems relies on solving a variety of optimization problems. Relevant applications include economic dispatch (ED), unit commitment (UC), market clearing, reserve planning, system restoration, expansion planning, and the optimal placement of distributed generation [1]. These optimization challenges encompass both transmission and distribution systems, with their complexity depending on factors such as problem type, decision time horizon, and the nature of the decision variables. Traditionally, grid optimization focused on a limited set of credible scenarios, where nonlinearity of power flow and integer variables mainly posed significant challenges. However, the integration of variable energy resources, flexible/controllable loads, and pronounced weather-related uncertainties have increased the complexity of grid operations significantly. Modern grids must now be optimized under a much broader range of scenarios, making optimization under uncertainty a far greater challenge that demands scalable and robust methods.
Optimization under uncertainty has been widely studied in the power systems domain, particularly in relation to various Optimal Power Flow (OPF)-based problems. In this context, the power systems community has explored a range of formulations, including stochastic, robust, distributionally robust, and chance-constrained optimization, as well as their hybrid forms [2]. Many real-world power system challenges, such as UC, ED, proactive dispatch, and investment planning, naturally lend themselves to a two-stage decision-making framework. In this setup, the first stage involves making non-anticipative decisions, while the second stage optimizes operational outcomes once uncertainty is realized. However, for large-scale systems, solving these two-stage problems becomes computationally intractable, particularly in resource-constrained environments [3]. This presents a critical gap in current methods, as the size and complexity of real-world power systems continue to grow. In this paper, we present a novel, computationally efficient, and scalable framework for solving two-stage stochastic optimization problems in power systems. The proposed method leverages a separable piecewise linear approximation of value function to achieve near-optimal solutions, significantly enhancing tractability for large-scale systems while maintaining high solution quality.
@@ I-A Background and Related Work
Ideally, two-stage stochastic optimization involves solving a deterministic extensive formulation (EF) where multiple copies of the second stage problem are added to the model for each scenario [4]. Solving such an extensive form is computationally intractable, as the size of the decision space increases exponentially with the number of scenarios. This is because the number of possible realizations increases exponentially with the number of uncertain parameters and the system size [5]. However, with the increasing level of uncertainty, it is not advisable to reduce excessively the number of scenarios just for computational purposes as a reasonable number of scenarios need to be considered in order to capture the uncertainty—otherwise, the obtained solutions may be sub-optimal or infeasible [6].
Traditional approaches to tackle scalability challenges include (1) solving the EF multiple times with reduced scenario size, e.g., using a Sample Average Approximation (SAA), (2) using a stage-wise decomposition, e.g., Benders decomposition (BD), and (3) using a scenario-wise decomposition, e.g., progressive hedging (PH) [3]. SAA involves solving the EF problem multiple times while taking a smaller sample size from the set of scenarios to approximate the value function. For example, SAA was used to minimize load consumption in distribution systems through conservation voltage reduction on the IEEE 37 bus system and 2000 scenarios [7]. However, for a realistic system size, the problem poses significant computational challenges even with a relatively small number of scenarios. BD and its variants have been utilized in solving various OPF problems. For instance, BD was used to solve the unit commitment problem for the IEEE 118 bus system considering 10 scenarios [8]. Stochastic decomposition, a variant of BD, has been applied in [9] for solving the economic dispatch problem considering a maximum system size of 96 buses and a maximum number of 100 scenarios. As the number of scenarios grows, the demand for generating valid cuts increases rapidly, making the problem intractable and unsuitable for larger systems with numerous scenarios. It was also observed that BD often leads to long computational times, excessive use of memory resources, poor feasibility and optimality cuts, and slow convergence (tail-off effect) [10]. On the other hand, PH, an application of the augmented Lagrangian relaxation, results in scenario subproblems which are then solved iteratively until convergence criteria are met [11]. PH was used in [12] for proactive dispatch and investment planning, considering the 2000 bus system and 512 scenarios. However, in real-world applications, the default settings for PH often pose challenges [13]. Moreover, due to its decompositional nature, PH is most effective in parallel architectures with multi-core devices, thereby restricting its broader applicability.
@@ I-B Contributions
We introduce Separable Projective Approximation Routine - Optimal Power Flow (SPAR-OPF), an efficient framework designed to solve two-stage stochastic optimization problems involving OPF in power systems. Leveraging a SPAR-based approach, our framework scales effectively with system size and can be implemented in resource-constrained environments. To the best of our knowledge, this is the first framework to solve OPF-based two-stage stochastic optimization problems using separable projective approximations of value function. From the application perspective, we consider the following two-stage setting: Stage-1 makes ex-ante decisions based on forecasts, and Stage-2 makes ex-post decisions after the realization of the event/uncertainty. Stage-1 can be thought of as a long-term or operational planning problem, and Stage-2 represents real-time operations. The general assumption is that planning decisions should be optimized considering optimal real-time operations. The computational performance of SPAR-OPF indicates its applicability for solving OPF under uncertainty for realistic power systems. The following are the main contributions of this paper.
@@ II-A Problem Formulation
We present SPAR-OPF as a general two-stage stochastic optimization model with optimal planning in the first stage and optimal operation (using OPF) in the second stage. Subsequently, specific details for the framework using DG siting and sizing as an example of a two-stage stochastic OPF problem are presented. The first stage problem is denoted as $\textbf{P}_{\textbf{g1}}$ and is defined in (1). It is important to mention that we do not impose any restrictions on the nature or structure of $\textbf{P}_{\textbf{g1}}$. The objective function consists of a first-stage cost $H(x,y)$ that is dependent only on the first-stage decision variables $x$ and $y$, and the expected recourse cost $Q(x,\xi)$. We obtain $Q(x,\xi)$ by solving the second stage OPF problem denoted by $\textbf{P}_{\textbf{g2}}$ with $z^{\xi}$ as decision variables, as given in (3). OPF has been shown to be exact under various convex relaxations and approximations [16]. We assume such convex problems in the second stage. Here, $x$ is the set of linking variables involved in the second stage for each scenario $\xi$. Note that the constraints in the second stage problem are separated into two types: one dependent on the first stage decision variable $x$, denoted as linking constraints (4), and the rest of the constraints (5)–(6) that do not directly involve the linking variables. Most general two-stage power grid optimization problems can be reformulated to fit the problem structure defined in (1)–(6).
[…]
@@
[…]
@@
If the second stage problem is convex with continuous recourse variables, $Q(x,\xi)$ is a piecewise linear convex function in the linking variable $x$.
@@
$\begin{gathered}L(x,\xi,z^{\xi},\alpha,\beta,\nu)=f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\end{gathered}$ (7)
The Lagrangian dual function $q(x,\xi,\alpha,\beta,\nu)$ and the associated dual problem optimal value $q^{D}(x,\xi)$ are then obtained as
[…]
$\begin{gathered}q(x,\xi,\alpha,\beta,\nu)=\inf_{z^{\xi}}~{}f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\quad\quad\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\quad\text{\rm and}\end{gathered}$ (8)
[…]
$\begin{gathered}q^{D}(x,\xi)=\max_{\alpha,\beta,\nu}~{}q(x,\xi,\alpha,\beta,\nu)\end{gathered}$ (9)
The Lagrangian dual problem (9) is convex. With the assumption that Slater’s condition is satisfied for the inequalities in (4) and (5), we get by strong duality that $Q(x,\xi)=q^{D}(x,\xi)$. From the expression for $q(x,\xi,\alpha,\beta,\nu)$ in (8), it follows that $q^{D}(x,\xi)$ is the maximum of affine functions of the linking variable $x$ through the term $-\alpha^{T}x$. Hence, $Q(x,\xi)$ is a piecewise linear convex function of $x$. ∎
Remark 1.
Under the assumption that the probability distribution of uncertain parameters is independent of decisions made and that Proposition 1 holds, $\mathbb{E_{\xi}}[Q(x,\xi)]$ is a piecewise linear convex function as it is the linear combination (coefficients are probabilities of scenarios) of piecewise linear convex functions.
II-B Value Function Approximation (VFA):
The value function $\mathbb{E_{\xi}}[Q(x,\xi)]$ in the objective function (1) of the first stage problem $\textbf{P}_{\textbf{g1}}$ is a function of linking variables $x$ and scenarios $\xi$ realized in the second stage. We write the value function as a summation of coordinate functions $g_{i}(x_{i})$ [^2] Coordinate function $g_{i}(x_{i})$ is a function of coordinate variable $x_{i}$ as given in (10) where $\mathcal{I}$ is the set of coordinate functions.
[…]
$G(x)=\mathbb{E}_{\xi}[Q(x,\xi)]=\sum_{i=1}^{|\mathcal{I}|}g_{i}(x_{i})$ (10)
Let us assume $g_{i}(x_{i})$ is a piecewise linear convex function with integer breakpoints in $x_{i}$ dimension as represented by the blue curve in Fig. 1. Let $m_{i}^{l}$ be its slope between points $l\ \in\mathcal{L}_{i}$ and $l+1$ as given in (11). The value of the function at any point $x_{i}$ can be obtained using (12) where $u$ is an integer point.
@@ II-B Value Function Approximation (VFA):
[…]
$\begin{gathered}g_{i}(x_{i})=\begin{cases}g_{i}(0)-\sum_{l=u}^{-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}<0\\ g_{i}(0)+\sum_{l=0}^{u-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}\geq 0\end{cases}\\ \text{where}\ u\leq x_{i}<u+1.\end{gathered}$ (12)
Figure 1: Piecewise linear convex functions with integer break points.
It is to be noted that, evaluating this function requires its value at origin, $g_{i}(0)$, which may not be available a priori. The decision points, however, are not affected by this constant term, and hence, $g_{i}(0)=0$ can be assumed. This shifts the function towards the origin as represented by the red curve in Fig. 1.
@@ II-B1 Lambda Method
[…]
@@ II-B2 Epigraph method
[…]
@@ II-C Separable Projective Approximation Routine (SPAR)
In this section, we present the SPAR algorithm from [15] for learning the coordinate functions $g_{i}$. Any term with the superscript $(k)$ refers to its value at iteration $k$. For simplicity, we represent $g_{i}(x_{i})$ as $g(x)$ and $m_{i}^{l}$ as $m_{l}$, which is the slope of function $g(x)$ between points $l$ and $l+1$. The overall goal is to learn and update the slopes iteratively such that the sequence $\{m^{(k)}\}$ converges to the true slope $\overline{m}$. In other words, the function $\{g^{(k)}\}$ constructed from the learned slopes using (12) is the functional approximation of the true function $g$. Algorithm 1 specifies how to learn the slopes of the coordinate function $g$. First, the slope vector $m^{(1)}$ is initialized to the zero vector and the maximum number of iterations is chosen as $k^{\max}$. Then, the first stage decision point $l^{(k)}$ is obtained by solving the first stage problem $\textbf{P}_{\textbf{g1}}$ after replacing the expected function by approximated function $G^{(k)}$ using (10). Notice that in initial iterations, exploration can be employed by randomly choosing decision points, rather than exploiting the optimal decisions in every iteration. After this, a random variable $\gamma^{(k)}$ is observed, which is $H_{k}$ measurable [^2] It means this value is based on the information available up to $k$ iterations, and its expectation is equal to the true slope $\overline{m}_{l^{(k)}}$. As discussed in [15], $\gamma^{(k)}$ satisfies this condition when Lagrange multipliers are taken after solving the second-stage problem for the given first-stage decision $l^{(k)}$. Next, the slope is updated at point $l^{(k)}$ as shown in step 5. Step size $\{\alpha^{(k)}\}$ is also $H_{k}$ measurable and assumed to satisfy $\sum_{k=1}^{\infty}\alpha^{(k)}=\infty$ and $\sum_{k=1}^{\infty}(\alpha^{(k)})^{2}<\infty$. The updated slope vector $n^{(k)}$ is then projected on a convex set $\mathcal{M}$ given by (19) to give slope vector $m^{(k+1)}$ for the next iteration to ensure the convexity of the learned function. Finally, the function $G^{(k)}$ is updated to give $G^{(k+1)}$ and the process continues until $k^{\max}$ iterations are reached or the termination criteria are met. One of the termination criteria considered in this paper is that the algorithm terminates when the moving average of the objective value at the current iteration falls within an $\epsilon$-radius of the moving average of the previous iteration.
Algorithm 1 Separable Projective Approximation Routine (SPAR)
@@
4:     Observe scenario and get random variable $\gamma^{(k)}$ after solving $\textbf{P}_{\textbf{g2}}$ such that,$\mathbb{E}(\gamma^{(k)}|l^{(1)},\dots,l^{(k)},\gamma^{(1)},...,\gamma^{(k-1)})=\overline{m}_{l^{(k)}}$

### v2

@@ I Introduction
The reliable, secure, and economical operation of power systems relies on solving a variety of optimization problems. Relevant applications include economic dispatch (ED), unit commitment (UC), market clearing, reserve planning, system restoration, expansion planning, and the optimal placement of distributed generation [1]. These optimization challenges encompass both transmission and distribution systems, with their complexity depending on factors such as problem type, decision time horizon, and the nature of the decision variables. Traditionally, grid optimization focused on a limited set of credible scenarios, where the nonlinearity of power flow and integer variables mainly posed significant challenges. However, the integration of variable energy resources, flexible/controllable loads, and pronounced weather-related uncertainties have increased the complexity of grid operations significantly [2]. Modern grids must now be optimized under a much broader range of scenarios, making optimization under uncertainty a far greater challenge that demands scalable and robust methods.
Optimization under uncertainty has been widely studied in the power systems domain, particularly in relation to various Optimal Power Flow (OPF)-based problems. In this context, the power systems community has explored a range of formulations, including stochastic, robust, distributionally robust, and chance-constrained optimization, as well as their hybrid forms [3]. Many real-world power system challenges, such as UC, ED, proactive dispatch, and investment planning, naturally lend themselves to a two-stage decision-making framework [4]. In this setup, the first stage involves making non-anticipative decisions, while the second stage optimizes operational outcomes once uncertainty is realized. However, for large-scale systems, solving these two-stage problems becomes computationally intractable, particularly in resource-constrained environments [5]. This presents a critical gap in current methods, as the size and complexity of real-world power systems continue to grow. In this paper, we present a novel, computationally efficient, and scalable framework for solving two-stage stochastic optimization problems in power systems. The proposed method leverages a separable piecewise linear approximation of value function to achieve near-optimal solutions, significantly enhancing tractability for large-scale systems while maintaining high solution quality.
@@ I-A Background and Related Work
Ideally, two-stage stochastic optimization involves solving a deterministic extensive formulation (EF) where multiple copies of the second stage problem are added to the model for each scenario [6]. Solving such an extensive form is computationally intractable, as the size of the decision space increases exponentially with the number of scenarios. With the increasing level of uncertainty, it is not advisable to excessively reduce the number of scenarios just for computational purposes, as a reasonable number of scenarios need to be considered in order to capture the uncertainty; otherwise, the obtained solutions may be sub-optimal or infeasible [2].
Traditional approaches to tackle scalability challenges include (1) solving the EF multiple times with reduced scenario size, e.g., using a Sample Average Approximation (SAA), (2) using a stage-wise decomposition, e.g., Benders decomposition (BD), and (3) using a scenario-wise decomposition, e.g., progressive hedging (PH) [5]. SAA involves solving the EF problem multiple times while taking a smaller sample size from the set of scenarios to approximate the value function. For example, SAA was used to solve the unit commitment problem considering the 118-bus system with a sample size of 50 scenarios [7]. However, for a realistic system size, the problem poses significant computational challenges even with a relatively small number of scenarios. BD and its variants have been utilized in solving various OPF problems. For instance, BD was used to solve the unit commitment problem for the IEEE 118 bus system considering 10 scenarios [8]. Stochastic decomposition, a variant of BD, has been applied in [9] for solving the economic dispatch problem considering a maximum system size of 96 buses and a maximum number of 100 scenarios. As the number of scenarios grows, the demand for generating valid cuts increases rapidly, making the problem intractable and unsuitable for larger systems with numerous scenarios. It was also observed that BD often leads to long computational times, excessive use of memory resources, poor feasibility and optimality cuts, and slow convergence (tail-off effect) [10]. PH has been applied to distributed generation (DG) siting and sizing on a 123-bus system with 400 scenarios [11], and to proactive dispatch and investment planning on a 2000-bus system with 512 scenarios [12]. However, in real-world applications, the default settings for PH often pose challenges [13]. Moreover, due to its decompositional nature, PH is most effective in parallel architectures with multi-core devices, thereby restricting its broader applicability.
@@ I-B Contributions
We introduce Separable Projective Approximation Routine - Optimal Power Flow (SPAR-OPF), an efficient framework designed to solve two-stage stochastic optimization problems involving OPF in power systems. Leveraging a SPAR-based approach, our framework scales effectively with system size and can be implemented in resource-constrained environments. To the best of our knowledge, this is the first framework to solve OPF-based two-stage stochastic optimization problems using separable projective approximation of value function. From the application perspective, we consider the following two-stage setting: Stage-1 makes ex-ante decisions based on forecasts, and Stage-2 makes ex-post decisions after the realization of the event/uncertainty. Stage-1 can be thought of as a long-term or operational planning problem, and Stage-2 represents real-time operations. The general assumption is that planning decisions should be optimized considering optimal real-time operations. The computational performance of SPAR-OPF indicates its applicability for solving OPF under uncertainty for realistic power systems. The following are the main contributions of this paper.
@@ II-A Problem Formulation
We present SPAR-OPF as a general two-stage stochastic optimization model with optimal planning in the first stage and optimal operation in the second stage. Subsequently, specific details for the framework using DG siting and sizing as an example of a two-stage stochastic OPF problem are presented. The first stage problem is denoted as $\textbf{P}_{\textbf{g1}}$ and is defined in (1). It is important to mention that we do not impose any restrictions on the nature or structure of $\textbf{P}_{\textbf{g1}}$. The objective function consists of a first-stage cost $H(x,y)$ that is dependent only on the first-stage decision variables $x$ and $y$, and the expected recourse cost $Q(x,\xi)$. We obtain $Q(x,\xi)$ by solving the second stage OPF problem denoted by $\textbf{P}_{\textbf{g2}}$ with $z^{\xi}$ as decision variables, as given in (3). OPF has been shown to be exact under various convex relaxations and approximations [16]. We assume such convex problems in the second stage with particular structures (see Remark 1) that the scientific community has widely considered when solving power system optimization problems under uncertainty [3, 6, 8, 17]. Here, $x$ is the set of linking variables involved in the second stage for each scenario $\xi$. Note that the constraints in the second stage problem are separated into two types: one dependent on the first stage decision variable $x$, denoted as linking constraints (4), and the rest of the constraints (5)–(6) that do not directly involve the linking variables. Most general two-stage power grid optimization problems can be reformulated to fit the problem structure defined in (1)–(6).
@@
 $\text{\quad s.t. }\quad A^{\xi}z^{\xi}\leq x\ \text{\{linking constraints\}}$ (4)
@@
If the second stage problem is convex with continuous recourse variables, $Q(x,\xi)$ is a convex function of the linking variable $x$.
@@
Let us consider $\alpha\geq 0$, $\beta\geq 0$ and $\nu$ be the dual variables corresponding to (4), (5) and (6) respectively. Note that $x$ is treated as parameter while solving second stage problem separately. Then, the Lagrangian of $\textbf{P}_{\textbf{g2}}$ is given as
$\begin{gathered}L(z^{\xi},\alpha,\beta,\nu;x,\xi)=f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\end{gathered}$ (7)
The Lagrangian dual function $q(\alpha,\beta,\nu;x,\xi)$ and the associated dual problem optimal value $q^{D}(x,\xi)$ are then obtained as
$\begin{gathered}q(\alpha,\beta,\nu;x,\xi)=\inf_{z^{\xi}}~f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\quad\quad\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\quad\text{\rm and}\end{gathered}$ (8)
$\begin{gathered}q^{D}(x,\xi)=\max_{\alpha,\beta,\nu}~q(\alpha,\beta,\nu;x,\xi)\end{gathered}$ (9)
The Lagrangian dual problem (9) is convex. With the assumption that Slater’s condition is satisfied for the inequalities in (4) and (5), which is a weak assumption in power systems problems [16], we get by strong duality that $Q(x,\xi)=q^{D}(x,\xi)$. Hence, $Q(x,\xi)$ is a convex function of $x$.
Corollary 1.
If the second stage problem is a linear problem (LP), $Q(x,\xi)$ is a piecewise linear convex function of the linking variable $x$.
Proof.
Immediately following the proof of Proposition 1, we see that $q(\alpha,\beta,\nu;x,\xi)$ is contained in a polyhedra in $(\alpha,\beta,\nu)$ due to second stage problem being LP. Therefore, $q^{D}(x,\xi)$ is the maximum of finitely many linear functions in $x$ and thus a piecewise linear convex function of $x$. Hence, $Q(x,\xi)$ is a piecewise linear convex function of $x$.
Lemma 1.
If the linking variable $x$ is integral, then there exists a piecewise linear function $F(x,\xi)$ such that $F(x,\xi)$ is equal to $Q(x,\xi)$ for all integer $x$. Consequently, the optimal solutions of $Q(x,\xi)$ over integer values of $x$ coincide with those of $F(x,\xi)$.
Proof.
Consider the function below.
$F(x,\xi)=Q(\lfloor x\rfloor,\xi)+\sum_{i=1}^{n}\left[Q(\lfloor x\rfloor+e_{i},\xi)-Q(\lfloor x\rfloor,\xi)\right]\left(x_{i}-\lfloor x_{i}\rfloor\right)$
where $\lfloor x\rfloor$ denotes the component-wise floor of $x$, $x_{i}$ is $i$-th component of $x$ and $e_{i}$ is the $i$-th unit vector.
Observe that $F$ is a piecewise linear extension of $Q$ with integer breakpoints. Trivially, $F(x,\xi)=Q(x,\xi)$ for all integer $x$ as desired. Furthermore, the optima of a piecewise linear function with integer break points occur at integer values, implying that the optima of the two functions coincide.
Remark 1.
We assume the function $Q(x,\xi)$ to be a piecewise linear convex function in the linking variable $x$. Corollary 1 ensures that this property holds for LP-based second-stage problems, which are widely used in two-stage stochastic power system problems [3]. Lemma 1 further extends the framework to integer linking variables with a general convex second-stage problem, including (but not limited to) system expansion planning [6], unit commitment [8], and DG siting and sizing [17]. For other convex formulations, Proposition 1 guarantees convexity in $x$ but not necessarily piecewise linearity; in such a case, $Q(x,\xi)$ may be approximated by a piecewise linear function so that the framework remains applicable.
Remark 2.
Under the common assumption of exogenous uncertainty [5], and that $Q(x,\xi)$ is piecewise linear convex in $x$, $\mathbb{E_{\xi}}[Q(x,\xi)]$ is a piecewise linear convex function as it is the linear combination, with probabilities of scenarios as coefficients, of piecewise linear convex functions.
II-B Value Function Approximation (VFA):
The value function $\mathbb{E_{\xi}}[Q(x,\xi)]$ in objective function (1) of first stage problem $\textbf{P}_{\textbf{g1}}$ is a function of linking variables $x$ and scenarios $\xi$ realized in the second stage. Two-stage problems, by nature, are not separable. However, separable approximations of value function have been widely used and found to provide near-optimal solution, typically by assuming separability with respect to linking variables when they appear affinely in the second-stage constraints  [14, 15]. We also consider such an approximation, and the value function is written as a summation of coordinate functions $g_{i}(x_{i})$ as given in (10) where $\mathcal{I}$ is the set of coordinate functions.
$G(x)=\mathbb{E}_{\xi}[Q(x,\xi)]=\sum_{i=1}^{|\mathcal{I}|}g_{i}(x_{i})$ (10)
@@ II-B Value Function Approximation (VFA):
$\hskip-8.67204pt\begin{gathered}g_{i}(x_{i})=\begin{cases}g_{i}(0)-\sum_{l=u}^{-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}<0\\ g_{i}(0)+\sum_{l=0}^{u-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}\geq 0\end{cases}\\ \text{where}\ u\leq x_{i}<u+1.\end{gathered}\hskip-8.67204pt$ (12)
It is to be noted that, evaluating this function requires its value at origin, $g_{i}(0)$, which may not be available a priori. The decision points, however, are not affected by this constant term, and hence, $g_{i}(0)=0$ can be assumed.
@@ II-B1 Lambda Method
 $\text{\hskip 43.36243pt s.t. }~~\sum_{l\in\mathcal{L}\,}\lambda_{i}^{l}=1~~\forall\ i\in\mathcal{I}\quad\quad$ (14)
@@ II-B2 Epigraph method
$\text{s.t. }\quad m_{i}^{l}\,(x_{i}-l)+g_{i}^{l}\leq t_{i}~\forall\ l\in\mathcal{L}_{i},\forall\ i\in\mathcal{I}$ (17)
@@ II-C Separable Projective Approximation Routine (SPAR)
In this section, we present the SPAR algorithm from [15] for learning the coordinate functions $g_{i}$. Any term with the superscript $(k)$ refers to its value at iteration $k$. For simplicity, we represent $g_{i}(x_{i})$ as $g(x)$ and $m_{i}^{l}$ as $m_{l}$, which is the slope of function $g(x)$ between points $l$ and $l+1$. The overall goal is to learn and update the slopes iteratively such that the sequence $\{m^{(k)}\}$ converges to the true slope $\overline{m}$. In other words, the function $\{g^{(k)}\}$ constructed from the learned slopes using (12) is the functional approximation of the true function $g$. Algorithm 1 specifies how to learn the slopes of the coordinate function $g$. First, the slope vector $m^{(1)}$ is initialized to the zero vector and the maximum number of iterations is chosen as $k^{\max}$. Then, the first stage decision point $l^{(k)}$ is obtained by solving the first stage problem $\textbf{P}_{\textbf{g1}}$ after replacing the expected function by approximated function $G^{(k)}$ using (10). Notice that in initial iterations, exploration can be employed by randomly choosing decision points, rather than exploiting the optimal decisions in every iteration. After this, a random variable $\gamma^{(k)}$ is observed, which is $H_{k}$ measurable, and its expectation is equal to the true slope $\overline{m}_{l^{(k)}}$. As discussed in [15], $\gamma^{(k)}$ satisfies this condition when Lagrange multipliers are taken after solving the second-stage problem for the given first-stage decision $l^{(k)}$. Next, the slope is updated at point $l^{(k)}$ as shown in step 5. Step size $\{\alpha^{(k)}\}$ is also $H_{k}$ measurable and assumed to satisfy $\sum_{k=1}^{\infty}\alpha^{(k)}=\infty$ and $\sum_{k=1}^{\infty}(\alpha^{(k)})^{2}<\infty$. The updated slope vector $n^{(k)}$ is then projected on a convex set $\mathcal{M}$ given by (20) to give slope vector $m^{(k+1)}$ for the next iteration to ensure the convexity of the learned function. Finally, the function $G^{(k)}$ is updated to give $G^{(k+1)}$ and the process continues until $k^{\max}$ iterations are reached or the termination criteria are met. One of the termination criteria considered in this paper is that the algorithm terminates when the moving average of the objective value at the current iteration falls within an $\epsilon$-radius of the moving average of the previous iteration.
The work in [15] had assumed relatively complete recourse, which means the second stage is always feasible for any first-stage decisions and slope can be updated with Lagrange multipliers. However, this assumption may not always hold. To make the framework more robust, we enhance SPAR algorithm by incorporating feasibility cuts that ensure the feasibility of the obtained decision by not allowing the search to go into infeasible directions. Specifically, if second stage problem is not feasible during an iteration, Farkas duals $\mu_{\zeta}$ corresponding to linking constraint (4) are obtained, and the feasibility cut of the form (18) is added to the first stage problem [19].
$\mu_{\zeta}^{T}x\geq 0$ (18)
Algorithm 1 Separable Projective Approximation Routine
@@
[…]